A polymer has the following molar mass distribution;

No. of molecules Molar Mass(g/mol)

50 5000

75 6000

Calculate:

i) Number ave. mol.wt

ii) Weight ave. mol. wt.

iii) Z- ave. mol.wt.

iv) PDI (12 marks)


1
Expert's answer
2021-02-15T03:00:21-0500

5000/50 = 100.

6000/75 = 80

11000/125 = 91,6666667

6000 x 1.2 = 7200.